8 years later I would be double that i was 5 years ago
Found 2 solutions by ikleyn, josgarithmetic:
Answer by ikleyn(52788)

.
x + 8 = 2*(x-5)


x + 8 = 2x - 10


8 + 10 = 2x - x


x = 18.

Solved.


Answer by josgarithmetic(39618)